OpenStudy (anonymous):

I'm not quite sure what to do with this sum \[\sum_{n=0}^{\infty} \frac{n!}{2*5*8*...*(3n+2)}\]

OpenStudy (experimentx):

are you trying to evaluate that sum or finding the bound for it??

OpenStudy (anonymous):

i'm supposed to determine if it's convergent or divergent

OpenStudy (anonymous):

does this mean \[ \frac{n!}{\infty (3n+2)}\]

OpenStudy (experimentx):

try using comparison test.

OpenStudy (anonymous):

sorry my above statement is incorrect

OpenStudy (anonymous):

the one with \[\infty\]

OpenStudy (anonymous):

would my comparison test be \[\sum_{n=0}^{\infty} \frac{n!}{(3n+2)}\]

OpenStudy (experimentx):

if you remove that 0 at the lower limit you will get \[ \frac{n!}{2*5*8*...*(3n+2)} \leq \frac{n!}{3*6*9*...*(3n)}\]

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} \frac{n!}{(3n+2)}\] would this do me any good?

OpenStudy (experimentx):

No wait \[ \sum_{n=0}^{\infty} \frac{n!}{2.5.8...(3n+2)} = 1/2 + \sum_{n=1}^{\infty} \frac{n!}{5.8...(3n+2)} \leq 1/2 + \sum_{n=1}^{\infty} \frac{n!}{3.6.9...(3n+2)}\]

OpenStudy (experimentx):

\[ \sum_{n=0}^{\infty} \frac{n!}{2.5.8...(3n+2)} = 1/2 + \sum_{n=1}^{\infty} \frac{n!}{5.8...(3n+2)} \leq 1/2 + \sum_{n=1}^{\infty} \frac{n!}{3.6.9...(3n)} \]

OpenStudy (experimentx):

\[ \sum_{n=1}^{\infty} \frac{n!}{3.6.9...(3n+2)} = \sum \frac{n!}{3^nn!} = \sum \frac 1{3^n }\] which converges by comparison test.

OpenStudy (anonymous):

how about the ratio test? or would that be silly?

OpenStudy (anonymous):

because that would eliminate the 2 5 8 etc

OpenStudy (experimentx):

no not silly .. try as you like ... but this is simpler this way.

OpenStudy (anonymous):

\[ \frac{(n+1)1}{2*5*8....etc}\ \frac{2*5*8....etc}{n!}\]

OpenStudy (anonymous):

so it's convergent

OpenStudy (anonymous):

\frac{(n+1)!}{2*5*8....etc}\ \frac{2*5*8....etc}{n!} sorry typing error

OpenStudy (anonymous):

\[\frac{(n+1)!}{2*5*8....etc}\ \frac{2*5*8....etc}{n!}\]

OpenStudy (experimentx):

yeah ... you will get  looks like ratio test is simpler.